OpenStudy (aliaaalihassan):

Applying The Comparison Test for Integral Convergence/Divergence, find if int: 1/x+e^x converges or diverges

OpenStudy (anonymous):

it is \[\int \frac{1}{x+e^x}dx\]?

OpenStudy (aliaaalihassan):

\[\int\limits_{0}^{\infty} (1)/(x+e^x)\]

OpenStudy (aliaaalihassan):

yes from 0 to infinity

OpenStudy (anonymous):

you could compare it to say \[\int \frac{dx}{e^x}\] which converges like crazy

OpenStudy (aliaaalihassan):

Can i also compare it to 1/x dx which also converges ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

since \[\int \frac{dx}{x}\] DIVERGES